Read the problem below and find the solution. Use a model or act the problem out to help solve it.
Holly's Bike Shop sells bicycles and tricycles. Holly wants to know how many bicycles she has in her store. She asks an employee to count for her. The employee counts them and says, "There is a total of 14 bicycles and tricycles in the store. The total number of wheels in the store is 34." How many bicycles are in the store?

Answer:?????

Answer by ankor@dixie-net.com(22740)

"There is a total of 14 bicycles and tricycles in the store.
The total number of wheels in the store is 34."
How many bicycles are in the store?
:
let b = no. of bikes
let t = no. trikes
:
Write an equation for each statement
"There is a total of 14 bicycles and tricycles in the store."
b + t = 14
t = (14-b), use this form for substitution
"The total number of wheels in the store is 34."
2b + 3t = 34
replace t with (14-b)
2b + 3(14-b) = 34
2b + 42 - 3b = 34
2b - 3b = 34 - 42
-b = -8
b = 8 bikes
:
:
Check this, find the no. of trikes
14 - 8 = 6 trikes
Check in the 2nd equation
2(8) + 3(6) = 34
16 + 18 = 34